OpenStudy (anonymous):

-6=-(x-5)-3(5+2x)-4(2x-4) solve for x.. i keep getting a negative non-integer which makes no sense, please help

OpenStudy (anonymous):

When faced with this kind of questions, if you are not confident of yourself, and prone to making careless mistakes, always do it step by step, don't need to rush it. -6=-(x-5)-3(5+2x)-4(2x-4) -6 = -x+5 -3(5+2x) - 4(2x-4) -6 = -x+5 -(15+6x) - (8x-16) -6 = -x+5 -15 -6x - 8x +16 -6 = -15x +6 15x = 12 x= 12/15 = 4/5

OpenStudy (anonymous):

hey wow i see my mistake thanks very much. i had -15x = 12, tyvm sir/maam!

OpenStudy (anonymous):

x=4/5

OpenStudy (anonymous):

thx np. Glad to be of help.